OpenStudy (anonymous):

A machine is set up to cut metal strips of varying lengths and widths based on the time (t) in minutes. The change in length is given by the function l(t)=t^2 - sq. root of t, and the change in width is given by w(t)=t^2 - 2t 1/2. Which function gives the change in area of the metal strips?

undeadknight26 (undeadknight26):

I would say...(I(t))(w(t))

OpenStudy (anonymous):

here the options to choose from a(t) = t^4+2t a(t)=t^4+2t+3t 5/2 a(t)=t^4-3t 5/2 +2t a (t)=t^4+2t-2t 1/2 + sq. rt. of t

undeadknight26 (undeadknight26):

In all honesty i didnt learn this yet..but I know that A = LW so we would have to combine the equations.

OpenStudy (anonymous):

That's a good start.

undeadknight26 (undeadknight26):

I honestly think this is C but not 100% sure... @ganeshie8 @shrutipande9 @waterineyes can you guys help please?

OpenStudy (anonymous):

what is w(t)??

OpenStudy (anonymous):

can you write it properly?

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

Is there any need of tagging?

OpenStudy (anonymous):

If yes, then tag entire OS members.. :P

undeadknight26 (undeadknight26):

Lol what do you mean mate? I didnt know if what i stated was right or wrong so i tagged people to give the person a better answer.

OpenStudy (anonymous):

\[Area = Length \times Width \implies A(t) = L(t) \cdot W(t)\]

OpenStudy (anonymous):

I am able to read l(t), but can anyone tell what is w(t)??

OpenStudy (anonymous):

\[l(t)=t^2 - \sqrt{t}\]

OpenStudy (anonymous):

\[w(t)=t^2 - 2 \cdot t^\frac{ 1}{2}\] ???

OpenStudy (anonymous):

@Shay96 If you have keyboard, then make use of it.. say something dear.. :P

OpenStudy (anonymous):

Sorry, on another question. But w(t) is t^2 - 2t 1/2

OpenStudy (anonymous):

What is 2t 1/2?

OpenStudy (anonymous):

1/2 is power ot t?

OpenStudy (anonymous):

1/2 is power of t?

OpenStudy (anonymous):

@undeadknight26 you are right.. :)

undeadknight26 (undeadknight26):

Awesome! *flex flex* im always right ;D